OpenStudy (anonymous):

use the distributive property to find the value of n : 15 x16=(10*10)+(10*n)+(5*10)+(5*n)

OpenStudy (anonymous):

240 = (100)+10n + (50) +5n 240 = 150 + 15n 90 = 15 n solve for n 90/15 = n n = 6

OpenStudy (anonymous):

are you sure thats the distributive way of solving.

OpenStudy (anonymous):

the distributive property allows you to multiply to get the product (what was inside the parentheses) and then add after. basically, you can do all the individual equations and then put them all together.

OpenStudy (anonymous):

also, note that I added the 10n to the 5n to get 15n, and the 100 to 50 to get 150. that's the distributive property, to put things that are alike together.

OpenStudy (anonymous):

oh pellet bro, i thought the distributive was multiplying an inside paranthesis numbers by an outside variable,!!

OpenStudy (anonymous):

but there wasn't any of that in the equation you put up.

OpenStudy (anonymous):

thats what threww me off, the funny thing i